# Moment maps and stability of holomorphic submersions

### v1

@@ 1. Introduction
In an appendix to this paper, we use the same approach to prove the following result on K-polystable deformations of cscK manifolds, already mentioned above. This fixes a mistake in Székelyhidi’s discussion [43, Proposition 8].
@@
The final step of the proof is to use the moment map flow for the finite-dimensional moment map; this is where the stability assumption is used. In particular, having a Kähler moment map is crucial for applying the moment map flow. Frequently used in the symplectic approach to GIT, the moment map flow is the flow along the gradient of the moment map squared; we mostly follow the formultation of Dervan–McCarthy–Sektnan [11, §4.2], which in turn builds on [22, 7], and we extend these works to apply to the case of non-trivial stabiliser. The flow starting from the point $w\in V_{\pi}^{+}$ that corresponds to a stable fibration $W\to B$ converges to a point $w_{\infty}$ inside $V_{\pi}^{+}$, representing a deformation of the fibration; this is where we use our “stability” result for the moment map flow. We prove that the stability of $W\to B$ implies that $w_{\infty}$ is a zero of the associated moment map and it belongs to the $K_{\pi,v}^{\mathbb{C}}$-orbit of $w$. Hence it is a solution to the optimal symplectic connection equation.
@@ Acknowledgments
I wish to thank Ruadhaí Dervan, Michael Hallam, Lars Martin Sektnan and Jacopo Stoppa for many discussions on the subject of this paper. I thank Andrés Ibáñez Núñez for discussions related to the moment map flow and Remark A.4. I thank the INdAM “National Group for Algebraic and Geometric Structures, and their Applications” (GNSAGA) for financial support. I thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme New equivariant methods in algebraic and differential geometry where work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1.
@@ 2. Preliminaries
Let $\mathrm{Aut}(M,L)$ be the group of automorphisms of $M$ which lift to $L$ and let $\mathfrak{h}_{0}$ be its Lie algebra. Then $\mathfrak{h}_{0}$ can be characterised as the space of holomorphic vector fields that can be written as the Riemannian gradient $\nabla_{g}f$, for some function $f$ called a holomorphy potential for the holomorphic vector field. Let $K=\mathrm{Isom}(M,\omega)$ be the group of holomorphic isometries of the Kähler metric $(\omega,J)$ and let $\mathfrak{k}$ be its Lie algebra. A well-known result of Matsushima and Lichnerowicz, known as the Matsushima criterion or the Cartan decomposition, states that when $\omega$ is cscK, the group $\mathrm{Aut}(M,L)$ is reductive [35, 33] (see also [21, §3.4]).
@@ 2.1. The moment map flow
It is clear from the definition of a Hamiltonian vector field that the Hamiltonian function is only unique up to a constant. The moment map then chooses a Hamiltonian function for the infinitesimal vector field.
@@
Let us view $V$ as an open ball inside a projective space $\mathbb{P}^{d}$: if $[x_{0}:\dots:x_{d}]$ are homogeneous coordinates on $\mathbb{P}^{d}$, the space $V$ can be viewed as an open ball around the origin in the affine open subset $x_{0}=0$. Following [11, Lemma 4.7], we can extend the action of $K$, the Kähler form $\omega$ and the moment map $\mu$ to $\mathbb{P}^{d}$ in such a way that their restriction to $V$ agrees with the original data. This allows us to appeal to results proven in the compact setting. Let $\mu_{\mathbb{P}^{d}}$ be the extended moment map. The moment map flow associated with $\mu_{\mathbb{P}^{d}}$ exists for all time. Moreover [29, Theorem 4.16] the limit set of the flow is a disjoint union of path-connected closed subsets where the norm square of the moment map has constant value. Let $C$ be one connected component and let $S_{C}$ be the stratum of points in $\mathbb{P}^{d}$ whose flow ends in $C$. A result of Duistermaat [32] guarantees that the flow starting at $x$ converges to a single point $x_{\infty}$ and the map
@@
is a deformation retraction. In particular the map
@@
is continuous. Let now $C$ be the connected component of the limit set that contains the origin of $V$; $C$ is open because the origin is a zero of the moment map, hence in particular is GIT-semistable by the Kempf–Ness theorem [28], and semistability is an open condition. Since $V$ is open there exists a neighbourhood $N^{\prime}(0)\subset C$ contained in $V$. From the continuity of the map (2.1), we can find a neighbourhood $N(0)$ in $V$ such that for each $x\in N(0)$ the limit point $x_{\infty}\in N^{\prime}(0)$. ∎
@@
Moreover, for the action of $K_{T^{\perp}}$, the point $x$ has now discrete stabiliser, and we can state the following result [11, Corollary 4.14].
@@ 2.2. K-stability
If $(\mathcal{X},\mathcal{L})$ is a test configuration, one can define the normalization of $(\mathcal{X},\mathcal{L})$ by taking the normalization $\widetilde{\mathcal{X}}$ of $\mathcal{X}$ and the pullback of $\mathcal{L}$ to $\widetilde{\mathcal{X}}$. The normalization is again a test configuration for $(X,L)$ [41, §5].
@@ 2.3. Deformation theory of Kähler metrics with constant scalar curvature
Let $\mathscr{G}$ be the group of Hamiltonian symplectomorphisms of $(M,\omega)$. Then $\mathscr{G}$ acts on $\mathscr{J}$ by pull-back. The Lie algebra of $\mathscr{G}$ can be identified with the space of smooth functions on $M$ with mean value zero, denoted $C_{0}^{\infty}(M)$.
@@
The deformations of the complex structure $J$ are parametrised by the finite-dimensional vector space
@@
The group $K$ can be complexified and from Matsushima’s criterion (Theorem 2.1) its complexification is $\mathrm{Aut}(M,L)$. In particular, we can realise the Lie algebra of $K$ as the space of real holomorphy potentials with respect to $(\omega,J)$. Changing the Kähler metric $\omega$ to $\omega+i\partial\bar{\partial}\varphi$ changes a holomorphy potential $h$ to
@@
see [44, Lemma 4.10]. In particular, it follows that the Lie algebra of $K$ can be identified with the kernel of $P$.
Given $\xi\in\mathfrak{k}$, we denote by $\sigma_{\xi}$ the corresponding infintesimal vector field and by $h_{\xi}$ the corresponding holomorphy potential. Moreover, we denote by
  1. (1)
    $\mathfrak{k}^{M}$ the realisation of $\mathfrak{k}$ as the space of holomorphy potentials with respect to $\omega$;
  2. (2)
The following theorem describes the deformations of the cscK complex structure $J$ and their scalar curvature. In the following, for an almost complex structure $J^{\prime}\in\mathscr{J}$, we denote by $\mathrm{Scal}(\omega,J^{\prime})$ the Hermitian scalar curvature of the Chern connection induced by $J^{\prime}$, which does not coincide with the Levi-Civita connection if $J^{\prime}$ is not integrable.
@@
    for every $x\in V$ there exists a Kähler potential $\varphi_{x}$ such that
    $\mathrm{Scal}(\omega+i\partial\bar{\partial}\varphi_{x},\Phi(x))-\widehat{S}\in\mathfrak{k}_{\varphi_{x}}^{M},$
@@
The first claim is mainly due to Kuranishi [30]; a proof adapted to take into account the compatibility with the symplectic form can be found in [7, §6]. We prove the second claim. For each $x\in V$, let $g_{x}$ be the Riemannian metric induced by $\omega$ and $\Phi(x)$. Consider the operator
@@
$(\xi,\varphi)$ $\mapsto\mathrm{Scal}(\omega+i\partial\bar{\partial}\varphi,\Phi(x))-\widehat{S}-h_{\xi}-\langle\nabla_{g_{x}}h_{\xi},\nabla_{g_{x}}\varphi\rangle_{g_{x}}.$
By our assumption, $G_{0}(0,0)=0$. The linearisation of $G_{x}$ at $(\xi,0)$ is the operator
$DG_{x}|_{(\xi,0)}(\varphi)=-\mathcal{D}^{*}_{x}\mathcal{D}_{x}(\varphi)+\langle\nabla_{g_{x}}\mathrm{Scal}(\omega,\Phi(x)),\nabla_{g_{x}}\varphi\rangle_{g_{x}}-h_{\xi}-\langle\nabla_{g_{x}}h_{\xi},\nabla_{g_{x}}\varphi\rangle_{g_{x}}$
Consider the operator
$L_{x}(\varphi)=-\mathcal{D}^{*}_{x}\mathcal{D}_{x}(\varphi)-h_{\xi}-\langle\nabla_{g_{x}}h_{\xi},\nabla_{g_{x}}\varphi\rangle_{g_{x}}.$
The kernel of $L_{x}$ is contained in $\ker\mathcal{D}^{*}_{x}\mathcal{D}_{x}$, which is contained in $\ker\mathcal{D}^{*}_{0}\mathcal{D}_{0}=\mathfrak{k}$, hence $L_{x}$ is surjective.
@@
$\lambda_{1}=\inf_{\psi\in\mathfrak{k}^{\perp}}\frac{\langle\Delta_{0}(\psi),\psi\rangle_{g_{0}}}{\lVert\psi\rVert^{2}_{g_{0}}},$
@@
¯$\lVert\bar{\partial}_{0}\psi\rVert^{2}_{L^{2}(g_{0})}\geq\lambda_{1}\lVert\psi\rVert^{2}_{L^{2}(g_{0})},\quad\psi\in\mathfrak{k}^{\perp}.$
@@
$\langle\Delta_{x}^{g_{0}}(\psi),\varphi\rangle_{g_{0}}=\lVert\bar{\partial}_{x}(\psi)\rVert^{2}_{g_{0}}+\lVert\bar{\partial}^{*}_{x}(\psi)\rVert^{2}_{g_{0}},$
@@
$\inf_{\psi\in\mathfrak{k}^{\perp}}\frac{\langle\Delta_{x}^{g_{0}}(\psi),\psi\rangle_{g_{0}}}{\lVert\psi\rVert^{2}_{g_{0}}}\geq\frac{\lambda_{1}}{2}.$
@@
¯$\lVert\bar{\partial}_{x}\psi\rVert^{2}_{L^{2}(g_{x})}\geq C\lVert\psi\rVert^{2}_{L^{2}(g_{x})},$
@@
$\lVert\mathcal{D}_{x}\psi\rVert^{2}_{L^{2}(g_{x})}\geq C\lVert\nabla_{g_{x}}\psi\rVert^{2}_{L^{2}(g_{x})}=C\lVert\bar{\partial}_{x}\psi\rVert^{2}_{L^{2}(g_{x})}\geq C\lVert\psi\rVert^{2}_{L^{2}(g_{x})}.$
Let now $(\xi,\psi)=Q_{x}(\varphi)$, where $\xi$ is the $L^{2}(g_{x})$-orthogonal projection of $\varphi$ on $\mathfrak{k}$. It follows that
$\lVert Q_{x}\varphi\rVert_{L^{2}(g_{x})}\leq\lVert\psi\rVert_{L^{2}(g_{x})}+\lVert\xi\rVert_{L^{2}(g_{x})}\leq C\lVert\varphi\rVert_{L^{2}(g_{x})}+\lVert\varphi\rVert_{L^{2}(g_{x})}\leq(C+1)\lVert\varphi\rVert_{L^{2}(g_{x})}.$
@@
$\lVert\psi\rVert_{W^{2,p+4}(g_{x})}\leq C\left(\lVert\psi\rVert_{L^{2}(g_{x})}+\lVert\varphi\rVert_{W^{2,p}(g_{x})}\right),$
@@
$\lVert Q_{x}\varphi\rVert_{W^{2,p}(g_{x})}\leq C\lVert\varphi\rVert_{W^{2,p}(g_{x})},$
where again the constants can be made independent on $x$. This proves the inequality (2.6).
Since $L_{x}$ differs from the acual linearisation $DG_{x}|_{(\xi,0)}$ by
$\langle\nabla_{g_{x}}\mathrm{Scal}(\omega,\Phi(x)),\nabla_{g_{x}}\psi\rangle_{g_{x}},$
and
$\mathrm{Scal}(\omega,\Phi(x))=\widehat{S}+O(\lvert x\rvert),$
the same estimates (2.6) holds for the right inverse to $DG_{x}|_{(\xi,0)}$, up to shrinking $V$. Therefore by the implicit function theorem for each $x$ we can find $(\xi,\varphi_{x})$, varying smoothly on $x$, such that $G_{x}(\xi,\varphi_{x})=0$. ∎
We will refer to $V$ as the Kuranishi space and to $\Phi$ as the Kuranishi map. Since we allow also non-integrable almost complex structures, the slice $V$ is an actual ball. Instead, in the original work by Kuranishi, the set $V$ parametrises only integrable complex structures, hence it is a complex analytic subspace of our $V$.

### v2

@@ 1. Introduction
In an appendix to this paper, we use the same approach to prove the following result on K-polystable deformations of cscK manifolds, already mentioned above. This fixes a mistake in Székelyhidi’s discussion [46, Proposition 8].
@@
The final step of the proof is to use the moment map flow for the finite-dimensional moment map; this is where the stability assumption is used. In particular, having a Kähler moment map is crucial for applying the moment map flow. Frequently used in the symplectic approach to GIT, the moment map flow is the flow along the gradient of the moment map squared; we mostly follow the formultation of Dervan–McCarthy–Sektnan [12, §4.2], which in turn builds on [24, 8], and we extend these works to apply to the case of non-trivial stabiliser. The flow starting from the point $w\in V_{\pi}^{+}$ that corresponds to a stable fibration $W\to B$ converges to a point $w_{\infty}$ inside $V_{\pi}^{+}$, representing a deformation of the fibration; this is where we use our “stability” result for the moment map flow. We prove that the stability of $W\to B$ implies that $w_{\infty}$ is a zero of the associated moment map and it belongs to the $K_{\pi,v}^{\mathbb{C}}$-orbit of $w$. Hence it is a solution to the optimal symplectic connection equation.
@@ Acknowledgments
I wish to thank Vestislav Apostolov, Ruadhaí Dervan, Michael Hallam, Lars Martin Sektnan, and Jacopo Stoppa for many discussions on the subject of this paper. I thank Andrés Ibáñez Núñez for discussions related to the moment map flow and Remark A.4. I thank the INdAM “National Group for Algebraic and Geometric Structures, and their Applications” (GNSAGA) for financial support. I thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme New equivariant methods in algebraic and differential geometry where work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1.
@@ 2. Preliminaries
Let $\mathrm{Aut}(M,L)$ be the group of automorphisms of $M$ which lift to $L$ and let $\mathfrak{h}_{0}$ be its Lie algebra. Then $\mathfrak{h}_{0}$ can be characterised as the space of holomorphic vector fields that can be written as the Riemannian gradient $\nabla_{g}f$, for some function $f$ called a holomorphy potential for the holomorphic vector field. Let $K=\mathrm{Isom}(M,\omega)$ be the group of holomorphic isometries of the Kähler metric $(\omega,J)$ and let $\mathfrak{k}$ be its Lie algebra. A well-known result of Matsushima and Lichnerowicz, known as the Matsushima criterion or the Cartan decomposition, states that when $\omega$ is cscK, the group $\mathrm{Aut}(M,L)$ is reductive [37, 35] (see also [23, §3.4]).
@@ 2.1. The moment map flow
It is clear from the definition of a Hamiltonian vector field that the Hamiltonian function is only unique up to a constant. The moment map then chooses a Hamiltonian function for the infinitesimal vector field. We next fix a bi-invariant inner product on $\mathfrak{k}$ with rational coefficients on the centre of $\mathfrak{k}$, and we use it to identify $\mathfrak{k}$ and $\mathfrak{k}^{*}$; therefore, we view the moment map as a map with values in $\mathfrak{k}$.
@@
Let us view $V$ as an open ball inside a projective space $\mathbb{P}^{d}$: if $[x_{0}:\dots:x_{d}]$ are homogeneous coordinates on $\mathbb{P}^{d}$, the space $V$ can be viewed as an open ball around the origin in the affine open subset $x_{0}=0$. Following [12, Lemma 4.7], we can extend the action of $K$, the Kähler form $\omega$, and the moment map $\mu$ to $\mathbb{P}^{d}$ in such a way that their restriction to $V$ agrees with the original data. This allows us to appeal to results proven in the compact setting. Let $\mu_{\mathbb{P}^{d}}$ be the extended moment map. The moment map flow associated with $\mu_{\mathbb{P}^{d}}$ exists for all time. Moreover [31, Theorem 4.16] the limit set of the flow is a disjoint union of path-connected closed subsets where the norm square of the moment map has constant value. Let $C$ be one connected component and let $S_{C}$ be the stratum of points in $\mathbb{P}^{d}$ whose flow ends in $C$. A result of Duistermaat [34] guarantees that the flow starting at $x\in S_{C}$ converges to a single point $x_{\infty}$ and the map
@@
is a deformation retraction. In particular, the map
@@
is continuous. Let now $C$ be the connected component of the limit set that contains the origin of $V$; $S_{C}$ is open because the origin is a zero of the moment map, hence in particular is GIT-semistable by the Kempf–Ness theorem [30], and semistability is an open condition. Since $V$ is open there exists a neighbourhood $N^{\prime}(0)\subset C$ contained in $V$. From the continuity of the map (2.1), we can find a neighbourhood $N(0)$ in $V$ such that for each $x\in N(0)$ the limit point $x_{\infty}\in N^{\prime}(0)$. ∎
@@
Moreover, for the action of $K_{T^{\perp}}$, the point $x$ has now discrete stabiliser, and we can state the following result [12, Corollary 4.14].
@@ 2.2. K-stability
If $(\mathcal{X},\mathcal{L})$ is a test configuration, one can define the normalization of $(\mathcal{X},\mathcal{L})$ by taking the normalization $\widetilde{\mathcal{X}}$ of $\mathcal{X}$ and the pullback of $\mathcal{L}$ to $\widetilde{\mathcal{X}}$. The normalization is again a test configuration for $(X,L)$ [44, §5].
@@ 2.3. Deformation theory of Kähler metrics with constant scalar curvature
Let $\mathscr{G}$ be the group of Hamiltonian symplectomorphisms of $(M,\omega)$. Then $\mathscr{G}$ acts on $\mathscr{J}$ by pull-back. The Lie algebra of $\mathscr{G}$ can be identified with the space of smooth functions on $M$ with mean value zero, denoted $C_{0}^{\infty}(M,\mathbb{R})$.
@@
The deformations of the complex structure $J$ which are integrable to first-order are parametrised by the finite-dimensional vector space
@@
The group $K$ can be complexified and from Matsushima’s criterion (Theorem 2.1) its complexification is $\mathrm{Aut}(M,L)$. In particular, we can realise the Lie algebra of $K$, denoted $\mathfrak{k}$, as the space of real holomorphy potentials with respect to $(\omega,J)$ with mean-value zero; it follows that the Lie algebra of $K$ can be identified with the kernel of $P$. Changing the Kähler metric $\omega$ to $\omega+i\partial\bar{\partial}\varphi$ changes a holomorphy potential $h$ to
@@
see [47, Lemma 4.10]. Given $\xi\in\mathfrak{k}$, we denote by $\sigma_{\xi}$ the corresponding infinitesimal vector field and by $h_{\xi}$ the corresponding holomorphy potential.
The following theorem describes the deformations of the cscK complex structure $J$ and their scalar curvature. In the following, when $J_{x}\in\mathscr{J}$ is a non-integrable almost complex structure, we denote by $g_{x}$ the Hermitian metric induced by $(\omega,J_{x})$ and by $\mathrm{Scal}(\omega,J_{x})$ its Chern-scalar curvature. Moreover, for a function $\varphi$ in the Sobolev space $W^{2,p}(g_{x})$ such that $\lVert\varphi\rVert_{W^{2,p}}\ll 1$, let
$\omega_{\varphi}:=\omega+\frac{1}{2}(\mathrm{d}\mathrm{d}^{c}\varphi)^{1,1},$
where the $(1,1)$-part is taken with respect to $J_{x}$ and $\mathrm{d}\mathrm{d}^{c}=-\mathrm{d}J_{x}\mathrm{d}$. While in general $\omega_{\varphi}$ is not closed, it satisfies the compatibility conditions
$\omega_{\varphi}(J_{x}\cdot,J_{x}\cdot)=\omega_{\varphi}(\cdot,\cdot),$
$\omega_{\varphi}(\cdot,J_{x}\cdot)>0,$
thus $(\omega_{\varphi},J_{x})$ induces a Hermitian metric. When $J_{x}$ is integrable, then $\omega_{\varphi}$ is closed and coincides with $\omega+i\partial\bar{\partial}\varphi$.
@@
    for every $x\in V$ there exists a smooth function $\varphi_{x}$ on $M$ and $\xi\in\mathfrak{k}$ such that
    $\mathrm{Scal}(\omega_{\varphi_{x}},\Phi(x))-\widehat{S}=h_{\xi}+\langle\nabla_{g_{x}}h_{\xi},\nabla_{g_{x}}\varphi_{x}\rangle_{g_{x}},$
@@
The first claim is mainly due to Kuranishi [32]; a proof adapted to take into account the compatibility with the symplectic form can be found in [8, §6]. We prove the second claim. For each $x\in V$, let $g_{x}$ be the Hermitian metric induced by $\omega$ and $\Phi(x)$. Consider the operator
@@
$(\xi,\varphi)$ $\mapsto\mathrm{Scal}(\omega_{\varphi},\Phi(x))-\widehat{S}-h_{\xi}-\langle\nabla_{g_{x}}h_{\xi},\nabla_{g_{x}}\varphi\rangle_{g_{x}}.$
By our assumption, $G_{0}(0,0)=0$. To compute the linearisation of $G_{x}$ at $(0,0)$ in the $\varphi$ variable we use the expression of the first variation of the Chern scalar curvature [1, Proposition 3.6] in the direction of $(\mathrm{d}\mathrm{d}^{c}\varphi)^{1,1}$. We first observe that, when $\mathrm{d}^{c}=-\Phi(x)\circ\mathrm{d}$, i.e. when we consider the almost complex structure $\Phi(x)$, then
$\mathrm{d}\mathrm{d}^{c}\varphi=(\mathrm{d}\mathrm{d}^{c}\varphi)^{1,1}+\frac{1}{2}\mathrm{d}\varphi\circ N_{x},$
where $N_{x}$ is the Nijenhuis tensor of $\Phi(x)$. Then the linearisation of $G_{x}$ is the operator [^1] More precisely, this is computed by applying [1, Proposition 3.6] to the variation given by $\left(-(\mathrm{d}\mathrm{d}^{c}\varphi)^{1,1}\right)^{\sharp_{\omega}}$, where $\sharp_{\omega}$ denotes the operation of raising an index using $\omega$.
$\begin{split}DG_{x}|_{(0,0)}(\varphi)=&-\Delta_{g_{x}}^{2}(\varphi)-\langle\mathrm{Ric}_{x},(\mathrm{d}\mathrm{d}^{c}\varphi)^{1,1}\rangle_{g_{x}}\\ &+\Delta_{g_{x}}\left(-\langle\mathrm{d}\varphi,\vartheta_{x}\rangle_{g_{x}}+\mathrm{Tr}_{g_{x}}\left(\frac{1}{2}\mathrm{d}\varphi\circ N_{x}\right)\right)+\langle\mathrm{Ric}_{x},\frac{1}{2}\mathrm{d}\varphi\circ N_{x}\rangle_{g_{x}}\\ &+\langle\mathrm{d}(-\Delta_{g}\varphi+\langle\mathrm{d}\varphi,\vartheta_{x}\rangle_{g_{x}})+\mathrm{d}\mathrm{Tr}_{g_{x}}\left(\frac{1}{2}\mathrm{d}\varphi\circ N_{x}\right),\vartheta_{x}\rangle_{g_{x}},\end{split}$
where $\vartheta_{x}$ is the torsion 1-form of the Chern connection of the metric $g_{x}$ and $\mathrm{Ric}_{x}$ is the Chern-Ricci form of $g_{x}$. Moreover,
$-\Delta_{g_{x}}^{2}(\varphi)-\langle\mathrm{Ric}_{x},(\mathrm{d}\mathrm{d}^{c}\varphi)^{1,1}\rangle_{g_{x}}=-\mathcal{D}^{*}_{x}\mathcal{D}_{x}(\varphi)+\langle\nabla_{g_{x}}\mathrm{Scal}(\omega,\Phi(x)),\nabla_{g_{x}}\varphi\rangle_{g_{x}},$
which is the linearisation of the scalar curvature when $(\omega,\Phi(x))$ is Kähler. Consider the Lichnerowicz operator
$L_{x}(\varphi)=-\mathcal{D}^{*}_{x}\mathcal{D}_{x}(\varphi).$
The kernel of $L_{x}$ is contained in $\ker\mathcal{D}^{*}_{0}\mathcal{D}_{0}=\mathfrak{k}$, hence $L_{x}$ is surjective.
@@
$\lambda_{1}=\inf_{\varphi\in\mathfrak{k}^{\perp}}\frac{\langle\Delta_{0}(\varphi),\varphi\rangle_{g_{0}}}{\lVert\varphi\rVert^{2}_{g_{0}}},$
@@
¯$\lVert\bar{\partial}_{0}\varphi\rVert^{2}_{L^{2}(g_{0})}\geq\lambda_{1}\lVert\varphi\rVert^{2}_{L^{2}(g_{0})},\quad\varphi\in\mathfrak{k}^{\perp}.$
@@
$\langle\Delta_{x}^{g_{0}}(\varphi),\varphi\rangle_{g_{0}}=\lVert\bar{\partial}_{x}(\varphi)\rVert^{2}_{g_{0}}+\lVert\bar{\partial}^{*}_{x}(\varphi)\rVert^{2}_{g_{0}},$
@@
$\inf_{\varphi\in\mathfrak{k}^{\perp}}\frac{\langle\Delta_{x}^{g_{0}}(\varphi),\varphi\rangle_{g_{0}}}{\lVert\varphi\rVert^{2}_{g_{0}}}\geq\frac{\lambda_{1}}{2}.$
@@
¯$\lVert\bar{\partial}_{x}\varphi\rVert^{2}_{L^{2}(g_{x})}\geq C\lVert\varphi\rVert^{2}_{L^{2}(g_{x})},$
@@
$\lVert\mathcal{D}_{x}\varphi\rVert^{2}_{L^{2}(g_{x})}\geq C\lVert\nabla_{g_{x}}\varphi\rVert^{2}_{L^{2}(g_{x})}=C\lVert\bar{\partial}_{x}\varphi\rVert^{2}_{L^{2}(g_{x})}\geq C\lVert\varphi\rVert^{2}_{L^{2}(g_{x})}.$
Let now $\varphi+\xi=Q_{x}(\psi)$, where $\xi$ is the $L^{2}(g_{x})$-orthogonal projection of $\psi$ on $\mathfrak{k}$. It follows that
$\lVert Q_{x}\psi\rVert_{L^{2}(g_{x})}\leq\lVert\varphi\rVert_{L^{2}(g_{x})}+\lVert\xi\rVert_{L^{2}(g_{x})}\leq C\lVert\psi\rVert_{L^{2}(g_{x})}+\lVert\psi\rVert_{L^{2}(g_{x})}\leq(C+1)\lVert\psi\rVert_{L^{2}(g_{x})}.$
@@
$\lVert\varphi\rVert_{W^{2,p+4}(g_{x})}\leq C\left(\lVert\psi\rVert_{L^{2}(g_{x})}+\lVert\psi\rVert_{W^{2,p}(g_{x})}\right),$
@@
$\lVert Q_{x}\psi\rVert_{W^{2,p}(g_{x})}\leq C\lVert\psi\rVert_{W^{2,p}(g_{x})},$
where again the constants can be made independent of $x$. This proves the inequality (2.6).
Since
$\begin{split}\mathrm{Scal}(\omega,\Phi(x))=\widehat{S}+O(\lvert x\rvert),\quad\vartheta_{x}=O(\lvert x\rvert),\quad N_{x}=O(\lvert x\rvert^{2}),\end{split}$
then $L_{x}(\varphi)$ differs from the acual linearisation $DG_{x}|_{(\xi,0)}(\varphi)$ by a term which is $O(\lvert x\rvert)$, so the same estimates (2.6) holds for the right inverse to $DG_{x}|_{(\xi,0)}$, up to shrinking $V$. Therefore by the implicit function theorem for each $x$ we can find $(\xi,\varphi_{x})$, varying smoothly on $x$, such that $G_{x}(\xi,\varphi_{x})=0$. ∎
We will refer to $V$ as the Kuranishi space and to $\Phi$ as the Kuranishi map. Since we also allow non-integrable almost complex structures, the slice $V$ is an actual ball. Instead, in the original work by Kuranishi, the set he constructs parametrises only integrable complex structures, hence it is a complex analytic subspace $V^{\mathrm{int}}$ of our $V$.